# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

### v2

@@ 2 Preliminaries
In this section, we recall definitions of semi-regular sequences and solving degrees, and collect some known facts related to them.
@@ 2.1 Semi-regular sequences
Definition 2.1.1 (Semi-regular sequences, [36, Definition 1]).
@@
When $K$ is an infinite field, Pardue also conjectured in [36, Conjecture B] that generic polynomial sequences are semi-regular.
We next review the notion of cryptographic semi-regular sequence, which is defined by a condition weaker than one for semi-regular sequence. The notion of cryptographic semi-regular sequence is introduced first by Bardet et al. (e.g., [2], [3]) motivated to analyze the complexity of computing Gröbner bases. Diem [14] also formulated cryptographic semi-regular sequences, in terms of commutative and homological algebra. The terminology “cryptographic” was named by Bigdeli et al. in their recent work [4], in order to distinguish such a sequence from a semi-regular one defined by Pardue (see Definition 2.1.1).
Definition 2.1.3 ([2, Definition 3]; see also [14, Definition 1]).
@@
Definition 2.1.7 ([2, Definition 4], [3, Definition 4]).
@@
As for an upper-bound on the degree of regularity, we refer to [23, Theorem 21].
@@
In Definition 2.1.7, since $R/I$ is Noetherian, it is Artinian if and only if it is of finite length. In this case, the degree of regularity $d_{\rm reg}(I)$ is equal to the Castelnuovo-Mumford regularity $\mathrm{reg}(I)$ of $I$ (see e.g., [16, §20.5] for the definition), whence $d_{\rm reg}(I)=\mathrm{reg}(I)=\mathrm{reg}(R/I)+1$.
Definition 2.1.9 ([2, Definition 5], [3, Definition 5]; see also [14, Section 2]).
@@
Proposition 2.1.10 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
@@
In 1985, Fröberg had already conjectured in [21] that, when $K$ is an infinite field, a generic sequence of homogeneous polynomials $f_{1},\ldots,f_{m}\in R$ of degrees $d_{1},\ldots,d_{m}$ generates an ideal $I$ with the Hilbert–Poincaré series of the form (2.1.2), namely $(f_{1},\ldots,f_{m})$ is cryptographic semi-regular. It can be proved (cf. [36]) that Fröberg’s conjecture is equivalent to Pardue’s one [36, Conjecture B]. We also note that Moreno-Socías conjecture [35] is stronger than the above two conjectures, see [36, Theorem 2] for a proof.
@@ 2.2 Solving degrees of Gröbner basis computation
In general, determining precisely the complexity of computing a Gröbner basis is very difficult; in the worst-case, the complexity is doubly exponential in the number of variables, see e.g., [10], [33], [37] for surveys. However, it is experimentally well-known that a Gröbner basis with respect to a graded monomial ordering, in particular degree reverse lexicographical (DRL) ordering, can be computed quite more efficiently than ones with respect to other orderings in general. Moreover, in the case where the input set $F=\{f_{1},\ldots,f_{m}\}$ of polynomials generate a zero-dimensional inhomogeneous ideal, once a Gröbner basis $G$ with respect to an efficient monomial ordering $\prec$ is computed, a Gröbner basis $G^{\prime}$ with respect to any other ordering $\prec^{\prime}$ can be computed easily by the FGLM basis conversion [20]. Even when $F$ is homogeneous, one can efficiently convert $G$ to $G^{\prime}$ by Gröbner walk [12] (or Hilbert driven [42] if both $\prec$ and $\prec^{\prime}$ are graded). From this, we focus on the case where the monomial ordering is graded, and if necessary we also assume that the ideal generated by the input polynomials is zero-dimensional.
@@
    As the first definition, we define the solving degree of an algorithm to compute a Gröbner basis as the highest degree of the polynomials involved during the execution of the algorithm, see [15, p. 36]. For example, applying Buchberger’s algorithm or its variants such as $F_{4}$ with the normal strategy, we collect critical S-pairs with the lowest degree and then reduce the corresponding S-polynomials in each iteration of the main loop of reductions. The lowest degree of each iteration is called the step degree. Then the solving degree is defined as the highest step degree. Instead, we may adopt the highest degree of $S$-polynomials appearing in the whole computation as in [41] and [40] by Semaeve-Tenti, and in this case we use the terminology “the solving degree in the strict sense”.
@@
    Their is a variant of the above first definition, where the solving degree is defined as a value depending not only on an algorithm but also on its implementation. More precisely, in [15, Section 2.1], the authors use the term solving degree for the step degree at which it takes the most amount of time among all iterations. In the cryptographic literature, the term solving degree often means this solving degree. Although this solving degree is estimated based on experiments, it is practically a quite important ingredient for analyzing the security of multivariate cryptosystems. The degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$ can be often used as a proxy for this solving degree.
@@
    We define the solving degree of $F$ with respect to a fixed (graded) monomial ordering as the lowest degree $d$ at which the reduced row echelon form (RREF) of $M_{\leq d}(F)$ produces a Gröbner basis of $F$.
@@
This strategy computes a basis of $V_{F,d}$ for each $d$, and therefore the smallest $d$ for which the mutant strategy terminates is equal to the solving degree of $F$ in terms of (III), see [24, Theorem 1]. As in [23], we refer to the algorithms such as Mutant-XL [7] and MXL2 [34] that employ this strategy as mutant algorithms. In the following, we denote the solving degree in (II) and that in (III) respectively by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$. By definitions, it is clear that $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq\mathrm{sd}_{\prec}^{\rm mac}(F)$ for any graded monomial oredering $\prec$, and the equality holds if the elements in $F$ are all homogeneous.
In a series of their celebrated works (cf. [8], [4], [24], [9], [23]), Gorla et al. provided a mathematical formulation for the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and algebraic invariants coming from $F$, such as the maximal Gröbner basis degree, the degree of regularity, the Castelnuovo–Mumford regularity, the first and last degrees, and so on. Here, the maximal Gröbner basis degree of the ideal $\langle F\rangle_{R}$ is the maximal degree of elements in the reduced Gröbner basis of $\langle F\rangle_{R}$ with respect to a fixed monomial ordering $\prec$, and is denoted by $\mathrm{max.GB.deg}_{\prec}(F)$. For any graded monomial oredering $\prec$, it is straightforward that
@@
In the non-homogeneous case, i.e., $F$ contains at least one non-homogeneous element, the euqalities in (2.2.1) do not hold in general, and it is not so easy to estimate any of the solving degrees. A straightforward way of bounding the solving degrees in the non-homogeneous case is to apply the homogenization as follows. We set $\prec$ as the DRL ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$, and fix it throughout the rest of this subsection. Let $y$ be an extra variable for homogenization as in Subsection A.2, and $\prec^{h}$ the homogenization of $\prec$, so that $y\prec x_{i}$ for any $i$ with $1\leq i\leq n$. Then, we have
@@
Theorem 2.2.1 (Lazard; [31, Theorem 2], [32, Théorèm 3.3]).
@@
With notation as above, let $d_{0}$ be a positive integer satisfying the following two properties:
@@
    The multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ is surjective.
@@
    For any $d\in\mathbb{Z}$ with $d\geq d_{0}$, the multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d+1}$ is injective.
Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$.
@@
Let $G$ be a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Clearly, we may suppose that each element of $G$ is homogeneous. It suffices to prove that $G_{\leq d_{0}}:=\{g\in G:\mathrm{deg}(g)\leq d_{0}\}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Indeed, the maximal degree of the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$ is not greater than that of any Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
Let $f\in\langle F^{h}\rangle$, and $d:=\mathrm{deg}(f)$. We show that there exists $g\in G_{\leq d_{0}}$ with $\mathrm{LM}(g)\mid\mathrm{LM}(f)$, by the induction on $d$. It suffices to consider the case where $f$ is homogeneous, since $\langle F^{h}\rangle$ is homogeneous. The case where $d\leq d_{0}$ is clear, and so we assume $d>d_{0}$.
First, if $\mathrm{LM}(f)\in R=K[x_{1},\ldots,x_{n}]$ (namely $y\nmid\mathrm{LM}(f)$), we choose an arbitrary monomial $t\in R$ of degree $d_{0}$ with $t\mid\mathrm{LM}(f)$. Since the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ by $y$ is surjective, there exists a homogeneous polynomial $h\in(R^{\prime})_{d_{0}-1}$ such that $h_{1}:=t-yh\in\langle F^{h}\rangle$. Here, $h_{1}$ is homogeneous of degree $d_{0}$, and $y\nmid t$, whence $\mathrm{LM}(h_{1})=t$. Therefore, we have $\mathrm{LT}(g)\mid t$ for some $g\in G$. Since $\mathrm{deg}(t)=d_{0}$, we also obtain $\mathrm{deg}(g)\leq d_{0}$, so that $g\in G_{\leq d_{0}}$.
Next, assume that $y\mid\mathrm{LM}(f)$. In this case, it follows from the definition of $\prec^{h}$ that any other term in $f$ is also divisible by $y$, so that $f\in\langle y\rangle$. Hence, we can write $f=yf_{1}$ for some homogeneous $f_{1}\in R^{\prime}$. By $d-1\geq d_{0}$, the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d}$ by $y$ is injective, so that $f_{1}\in\langle F^{h}\rangle_{d-1}$. By the induction hypothesis, there exists $g\in G_{\leq d_{0}}$ such that $\mathrm{LM}(g)\mid\mathrm{LM}(f_{1})$. Since $\mathrm{LM}(f)=y\mathrm{LM}(f_{1})$, we obtain $\mathrm{LM}(g)\mid\mathrm{LM}(f)$. We have proved that $G_{\leq d_{0}}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. ∎
Lazard proved that we can take $d_{1}+\cdots+d_{\ell}-\ell+1$ in Theorem 2.2.1 as $d_{0}$ in Lemma 2.2.2. Lazard’s bound given in (2.2.2) is also referred to as the Macaulay bound, and it provides an upper-bound for the solving degree of $F$ with respect to a DRL ordering.
@@
for any graded monomial ordering $\prec^{\prime}$ on $R$, see [8, Remark 15] or Lemma 4.2.4 below for a proof. Both $d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{\rm sd}_{\prec}^{\rm mac}(F)$ are greater than or equal to ${\rm max.GB.deg}_{\prec}(F)$, whereas it is pointed out in [4], [8], and [9] by explicit examples that any of the degree of regularity and the first fall degree does not produce an estimate for the solving degrees in general, even when $F$ is an affine (cryptographic) semi-regular sequence. Caminata-Gorla proved in [9] that the solving degree $\mathrm{sd}_{\prec}^{\rm mut}(F)$ is nothing but the last fall degree if it is greater than the maximal Gröbner basis degree:
Theorem 2.2.3 ([9, Theorem 3.1]).
With notation as above, for any degree-compatible monomial ordering $\prec^{\prime}$ on $R$, we have the following equality:
$\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\color[rgb]{0,0,0}\prec^{\prime}}}(F)\},$
where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]).
@@
Theorem 2.2.5 ([40, Theorem 2.1], [41, Theorem 3.65 & Corollary 3.67]).

### v3

@@ 2 Preliminaries
In this section, we recall definitions of semi-regular sequences and solving degrees, and collect some known facts related to them. Subsequently, we also extend the notion of degree of regularity.
@@ 2.1 Semi-regular sequences
Definition 2.1.1 (Semi-regular sequences, [37, Definition 1]).
@@
When $K$ is an infinite field, Pardue also conjectured in [37, Conjecture B] that generic polynomial sequences are semi-regular.
We next review the notion of cryptographic semi-regular sequence, which is defined by a condition weaker than one for semi-regular sequence. The notion of cryptographic semi-regular sequence is introduced first by Bardet et al. (e.g., [2], [4]) motivated to analyze the complexity of computing Gröbner bases. Diem [15] also formulated cryptographic semi-regular sequences, in terms of commutative and homological algebra. The terminology ‘cryptographic’ was named by Bigdeli et al. in their recent work [5], in order to distinguish such a sequence from a semi-regular one defined by Pardue (see Definition 2.1.1).
Definition 2.1.3 ([2, Definition 3]; see also [15, Definition 1]).
@@
Definition 2.1.7 ([2, Definition 4], [4, Definition 4]).
@@
As for an upper-bound on the degree of regularity, we refer to [24, Theorem 21]. An elementary but important fact that relates $d_{\rm reg}(I)$ and ${\rm max.GB.deg}_{\prec}(I)$ is the following (the proof is straightforward, but write it here for the readers’ convenience):
Lemma 2.1.8.
For any homogeneous ideal $I$ of $R$ and any graded ordering $\prec$ on the set of monomials in $R$, we have ${\rm max.GB.deg}_{\prec}(I)\leq d_{\rm reg}(I)$.
Proof.
The case $d_{\rm reg}(I)=\infty$ is trivial, so we consider the case where $d_{\rm reg}(I)<\infty$, namely $R/I$ is Artinian. Let $G$ be the reduced Gröbner basis of $I$ with respect to $\prec$, and put $D:=d_{\rm reg}(I)$. Assume for a contradiction that there were an element $g\in G$ such that $\deg(g)>D$. Since $\prec$ is graded, we have $\deg(g)=\deg\mathrm{LM}(g)$, so we choose any monomial $M\in R_{D}$ dividing $\mathrm{LM}(g)$. Then, it follows from $R_{D}=I_{D}$ that $M$ is divisible by $\mathrm{LM}(g^{\prime})$ for some $g^{\prime}\in G$ with $\deg(g^{\prime})=\deg\mathrm{LM}(g^{\prime})\leq D$, so that $\mathrm{LM}(g)$ is divisible by $\mathrm{LM}(g^{\prime})$ with $g\neq g^{\prime}$. This contradicts to that $G$ is reduced. ∎
@@
In Definition 2.1.7, since $R/I$ is Noetherian, it is Artinian if and only if it is of finite length. In this case, the degree of regularity $d_{\rm reg}(I)$ is equal to the Castelnuovo-Mumford regularity $\mathrm{reg}(I)$ of $I$ (see e.g., [17, §20.5] for the definition), whence $d_{\rm reg}(I)=\mathrm{reg}(I)=\mathrm{reg}(R/I)+1$.
Definition 2.1.10 ([2, Definition 5], [4, Definition 5]; see also [15, Section 2]).
@@
Proposition 2.1.11 ([15, Proposition 1 (d)]; see also [4, Proposition 6]).
@@
In 1985, Fröberg had already conjectured in [22] that, when $K$ is an infinite field, a generic sequence of homogeneous polynomials $f_{1},\ldots,f_{m}\in R$ of degrees $d_{1},\ldots,d_{m}$ generates an ideal $I$ with the Hilbert–Poincaré series of the form (2.1.2), namely $(f_{1},\ldots,f_{m})$ is cryptographic semi-regular. It can be proved (cf. [37]) that Fröberg’s conjecture is equivalent to Pardue’s one [37, Conjecture B]. We also note that Moreno-Socías conjecture [36] is stronger than the above two conjectures, see [37, Theorem 2] for a proof.
@@ 2.2 Solving degrees of Gröbner basis computation
In general, determining precisely the complexity of computing a Gröbner basis is very difficult; in the worst-case, the complexity is doubly exponential in the number of variables, see e.g., [11], [34], [38] for surveys. However, it is experimentally well-known that a Gröbner basis with respect to a graded monomial ordering, in particular degree reverse lexicographical (DRL) ordering, can be computed quite more efficiently than ones with respect to other orderings in general. Moreover, in the case where the input set $F=\{f_{1},\ldots,f_{m}\}$ of polynomials generate a zero-dimensional inhomogeneous ideal, once a Gröbner basis $G$ with respect to an efficient monomial ordering $\prec$ is computed, a Gröbner basis $G^{\prime}$ with respect to any other ordering $\prec^{\prime}$ can be computed easily by the FGLM basis conversion [21]. Even when $F$ is homogeneous, one can efficiently convert $G$ to $G^{\prime}$ by Gröbner walk [13] (or Hilbert driven [43] if both $\prec$ and $\prec^{\prime}$ are graded). From this, we focus on the case where the monomial ordering is graded, and if necessary we also assume that the ideal generated by the input polynomials is zero-dimensional (see Terminology in Section 1 for the meaning of zero-dimensional).
@@
    As the first definition, we define the solving degree of an algorithm to compute a Gröbner basis as the highest degree of the polynomials involved during the execution of the algorithm, see [16, p. 36]. For example, applying Buchberger’s algorithm or its variants such as $F_{4}$ with the normal strategy (or called normal selection strategy, see [14, \$II.10]), we collect critical S-pairs with the lowest degree and then reduce the corresponding S-polynomials in each iteration of the main loop of reductions. The lowest degree of each iteration is called the step degree. Then the solving degree is defined as the highest step degree. Instead, we may adopt the highest degree of $S$-polynomials appearing in the whole computation as in [42] and [41] by Semaeve-Tenti, and in this case we use the terminology ‘the solving degree in the strict sense’.
@@
    Their is a variant of the above first definition, where the solving degree is defined as a value depending not only on an algorithm but also on its implementation. More precisely, in [16, Section 2.1], the authors use the term solving degree for the step degree at which it takes the most amount of time among all iterations. In the cryptographic literature, the term solving degree often means this solving degree. Although this solving degree is estimated based on experiments, it is practically a quite important ingredient for analyzing the security of multivariate cryptosystems. The degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$ can be often used as a proxy for this solving degree.
@@
    We define the solving degree of $F$ with respect to a fixed (graded) monomial ordering $\prec$ as the lowest degree $d$ at which the reduced row echelon form (RREF) of $M_{\leq d}(F)$ produces a Gröbner basis of $F$ with respect to $\prec$.
@@
This strategy computes a basis of $V_{F,d}$ for each $d$, and therefore the smallest $d$ for which the mutant strategy terminates is equal to the solving degree of $F$ in terms of (III), see [25, Theorem 1]. As in [24], we refer to the algorithms such as Mutant-XL [8] and MXL2 [35] that employ this strategy as mutant algorithms. In the following, we denote the solving degree in (II) and that in (III) respectively by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$. By definitions, it is clear that $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq\mathrm{sd}_{\prec}^{\rm mac}(F)$ for any graded monomial oredering $\prec$, and the equality holds if the elements in $F$ are all homogeneous.
In a series of their celebrated works (cf. [9], [5], [25], [10], [24]), Gorla et al. provided a mathematical formulation for the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and algebraic invariants coming from $F$, such as the maximal Gröbner basis degree, the degree of regularity, the Castelnuovo–Mumford regularity, the first and last fall degrees, and so on. Here, the maximal Gröbner basis degree of the ideal $\langle F\rangle_{R}$ is the maximal degree of elements in the reduced Gröbner basis of $\langle F\rangle_{R}$ with respect to a fixed monomial ordering $\prec$, and is denoted by $\mathrm{max.GB.deg}_{\prec}(F)$. For any graded monomial oredering $\prec$, it is straightforward that
@@
In the inhomogeneous case, i.e., $F$ contains at least one inhomogeneous element, the equalities in (2.2.1) do not hold in general, and it is not so easy to estimate any of the solving degrees. A straightforward way of bounding the solving degrees in the inhomogeneous case is to apply the homogenization as follows. We set $\prec$ as the DRL ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$, and fix it throughout the rest of this subsection. Let $y$ be an extra variable for homogenization as in Subsection A.2, and $\prec^{h}$ the homogenization of $\prec$, so that $y\prec x_{i}$ for any $i$ with $1\leq i\leq n$. Then, we have
@@
Theorem 2.2.1 (Lazard; [32, Theorem 2], [33, Théorèm 3.3]).
@@
Let $I$ be a homogeneous ideal of $R^{\prime}$, and let $d_{0}$ be a positive integer satisfying the following two properties:
@@
    The multiplication-by-$y$ map $(R^{\prime}/I)_{d_{0}-1}\longrightarrow(R^{\prime}/I)_{d_{0}}$ is surjective.
@@
    For any $d\in\mathbb{Z}$ with $d\geq d_{0}$, the multiplication-by-$y$ map $(R^{\prime}/I)_{d}\longrightarrow(R^{\prime}/I)_{d+1}$ is injective.
Then we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(I)\leq d_{0}$, where $\prec^{\prime}$ is a homogenization (with respect to $y$) of any graded monomial ordering on $R$, see Subsection A.2 below for details.
@@
Let $G$ be a Gröbner basis of $I$ with respect to $\prec^{\prime}$. Clearly, we may suppose that each element of $G$ is homogeneous. It suffices to prove that $G_{\leq d_{0}}:=\{g\in G:\mathrm{deg}(g)\leq d_{0}\}$ is a Gröbner basis of $I$ with respect to $\prec^{\prime}$. Indeed, the maximal degree of the reduced Gröbner basis of $I$ with respect to $\prec^{\prime}$ is not greater than that of any Gröbner basis of $I$ with respect to $\prec^{\prime}$.
Let $f\in I$, and $d:=\mathrm{deg}(f)$. We show that there exists $g\in G_{\leq d_{0}}$ with $\mathrm{LM}(g)\mid\mathrm{LM}(f)$, by the induction on $d$. It suffices to consider the case where $f$ is homogeneous, since $I$ is homogeneous. The case where $d\leq d_{0}$ is clear, and so we assume $d>d_{0}$.
First, we consider the case where $y\nmid\mathrm{LM}(f)$ (namely $\mathrm{LM}(f)\in R=K[x_{1},\ldots,x_{n}]$). We choose an arbitrary monomial $t\in R$ of degree $d_{0}$ with $t\mid\mathrm{LM}(f)$. Since the multiplication map $(R^{\prime}/I)_{d_{0}-1}\longrightarrow(R^{\prime}/I)_{d_{0}}$ by $y$ is surjective, there exists a homogeneous polynomial $h\in(R^{\prime})_{d_{0}-1}$ such that $h_{1}:=t-yh\in I$. Here, $h_{1}$ is homogeneous of degree $d_{0}$, and $y\nmid t$, whence $\mathrm{LM}(h_{1})=t$. Therefore, we have $\mathrm{LT}(g)\mid t$ for some $g\in G$. Since $\mathrm{deg}(t)=d_{0}$, we also obtain $\mathrm{deg}(g)\leq d_{0}$, so that $g\in G_{\leq d_{0}}$.
Next, assume that $y\mid\mathrm{LM}(f)$. In this case, it follows from the definition of $\prec^{\prime}$ that any other term in $f$ is also divisible by $y$, so that $f\in\langle y\rangle$. Hence, we can write $f=yf_{1}$ for some homogeneous $f_{1}\in R^{\prime}$. By $d-1\geq d_{0}$, the multiplication map $(R^{\prime}/I)_{d-1}\longrightarrow(R^{\prime}/I)_{d}$ by $y$ is injective, so that $f_{1}\in I_{d-1}$. By the induction hypothesis, there exists $g\in G_{\leq d_{0}}$ such that $\mathrm{LM}(g)\mid\mathrm{LM}(f_{1})$. Since $\mathrm{LM}(f)=y\mathrm{LM}(f_{1})$, we obtain $\mathrm{LM}(g)\mid\mathrm{LM}(f)$. We have proved that $G_{\leq d_{0}}$ is a Gröbner basis of $I$ with respect to $\prec^{\prime}$. ∎
Lazard proved that we can take $d_{1}+\cdots+d_{\ell}-\ell+1$ in Theorem 2.2.1 as $d_{0}$ in Lemma 2.2.2, where we also take $I$ and $\prec^{\prime}$ to be $\langle F^{h}\rangle$ and $\prec^{h}$ respectively. Lazard’s bound given in (2.2.2) is also referred to as the Macaulay bound, and it provides an upper-bound for the solving degree of $F$ with respect to a DRL ordering.
@@
for any graded monomial ordering $\prec^{\prime}$ on $R$, see [9, Remark 15] or Lemma 4.2.4 below for a proof. Both $d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{\rm sd}_{\prec}^{\rm mac}(F)$ are greater than or equal to ${\rm max.GB.deg}_{\prec}(F)$, whereas it is pointed out in [5], [9], and [10] by explicit examples that any of the degree of regularity and the first fall degree does not produce an estimate for the solving degrees in general, even when $\bm{F}$ is an affine (cryptographic) semi-regular sequence. Caminata-Gorla proved in [10] that the solving degree $\mathrm{sd}_{\prec}^{\rm mut}(F)$ is nothing but the last fall degree $d_{F}$ if $d_{F}$ is greater than the maximal Gröbner basis degree:
Theorem 2.2.3 ([10, Theorem 3.1]).
With notation as above, for any graded monomial ordering $\prec^{\prime}$ on $R$, we have the following equality:
$\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{\prec^{\prime}}(F)\},$
where $d_{F}$ denotes the last fall degree of $F$ defined in [10, Definition 1.5] (originally in [28], [27]).
@@
Theorem 2.2.5 ([41, Theorem 2.1], [42, Theorem 3.65 & Corollary 3.67]).